# Robust Semantic Communications for Speech Transmission

### v2

@@ Abstract
In this paper, we propose a robust semantic communication system for speech transmission, named Ross-S2T, by delivering the essential semantic information. Particularly, we consider the speech-to-text translation (S2TT) as the transmission goal. First, a deep semantic encoder is developed to directly convert speech in the source language to textual features associated with the target language, facilitating the end-to-end (E2E) semantic exchange to perform the S2TT task and reducing the transmission data without performance degradation. To mitigate semantic impairments inherent in the corrupted speech, a novel generative adversarial network (GAN)-enabled deep semantic compensator is established to estimate the lost semantic information within the speech and extract deep semantic features simultaneously, which enables robust semantic transmission for corrupted speech. Furthermore, a semantic probe-aided compensator is devised to enhance the semantic fidelity of recovered semantic features and improve the understandability of the target text. According to simulation results, the proposed Ross-S2T exhibits superior S2TT performance compared to conventional approaches and high robustness against semantic impairments.
@@ I Introduction
According to the three-level communication architecture proposed by Shannon and Weaver [3], semantic communications are the second level of communications that prioritize conveying the underlying meaning by representing the input message with minimal ambiguity through semantics, which overcomes the limitation of conventional communications to process data at the bit level and drives the evolution of intelligent communications. However, the investigation of semantic communication is in its infancy and there is no consensus on the definition of semantics, hindering the motivation to represent the semantic information with a rigorous formula. Inspired by the thriving of artificial intelligence (AI) in diverse areas, deep learning (DL)-enabled semantic communication paradigm breaks the bottleneck of a mathematical theory to quantify the semantics and has shown its great potential to learn semantic information by designing sophisticated neural networks (NNs). DL-enabled semantic communications have experienced unprecedented developments due to the ubiquity of intelligent mobile devices and the booming demand for semantic-driven data transmission.
@@
According to the transmission goal, semantic communication systems can restore the source message and/or perform the AI tasks. The global semantic features are transmitted to achieve the data reconstruction. However, in the task-oriented semantic transmission mechanism, only the task-related semantic features are extracted, serving diverse user requests and enabling efficient transformation between multimodal data. Particularly, the pioneering work on DL-enabled semantic communications, named DeepSC [4], has been proposed to recover accurate text by leveraging the transformer module to learn textual semantic features. A variant of DeepSC, named R-DeepSC [5], has been devised to eliminate semantic noise and facilitate robust text transmission. In [6], Weng et al. introduced a semantic communication system for speech transmission, named DeepSC-S, which utilizes a joint semantic-channel coding scheme to extract and transmit global semantic features. Inspired by DeepSC-S, a deep speech semantic transmission scheme has been developed in [7] by adopting a flexible rate-distortion trade-off to achieve end-to-end (E2E) optimization. Huang et al. [8] considered a semantic communication system for image transmission and Jiang et al. [9] presented a video semantic transmission paradigm to alleviate semantic errors by incorporating a semantic detector.
@@
Furthermore, Xie et al. [10] established a task-oriented semantic communications system for machine translation and visual question-answering tasks by fusing textual and visual semantic features. The speech recognition and speech synthesis tasks are performed in an efficient speech semantic transmission scheme by converting the speech input into the task-related semantic features [11]. In [12], Xu et al. proposed reinforcement learning-enabled semantic communications for scene classification in unmanned aerial systems. Zhang et al. investigated a semantic communication system for extended reality (XR) tasks by transmitting highly compressed semantic information to reduce network traffic. In [13], A unified multimodal multi-task semantic communication architecture, named U-DeepSC, has been developed by sharing trainable parameters amongst various tasks to reduce the redundancy of semantic features and accelerate the inference process.
@@
In this paper, a robust semantic communication system for speech transmission, named Ross-S2T, is proposed. We argue that existing research works on task-oriented semantic communications for speech only extract textual semantic features constrained to the source language, i.e., shallow semantic features, encouraging the exploration of deep semantic features spanning various languages. Moreover, the intractable semantic impairments inherent in the corrupted speech are investigated. In this context, the speech-to-text translation (S2TT) task is considered in semantic transmission scenarios with corrupted speech input. The contributions of this paper are summarized as follows:
  • A semantic communication system for S2TT in the context of clear speech, named DeepSC-S2T, is developed by utilizing a deep semantic encoder to extract textual semantic features related to the target language from speech in the source language.
@@
  • According to our comprehensive literature review, the speech semantic impairments in semantic communications are not investigated. We propose a generative adversarial network (GAN)-enabled deep semantic compensator to estimate the damaged semantic information inherent in the corrupted speech and generate deep semantic features simultaneously.
@@
The rest of this paper is structured as follows. In Section II, the model of the robust semantic communications for S2TT is provided. Section III introduces the details of the proposed Ross-S2T. Section IV presents experimental results and Section V conclusions this paper.
@@ II System Model
In this section, we introduce robust semantic communication systems for speech transmission and consider S2TT as the transmission goal. The considered system aims to address two primary challenges. The first challenge is to deliver E2E semantic exchange and achieve efficient transmission from speech in the source language to text in the target language. The second one is to devise a semantic impairment suppression mechanism to contend with semantic impairments within the corrupted speech. To this end, the novel deep semantic codec mechanism is established to facilitate speech transmission for S2TT, and the deep semantic compensator is first developed to compensate for the complicated semantic impairments.
Figure 1: The model structure of robust semantic communications for speech-to-text translation.
@@ II-A Clear Speech Input
The designed system is tailored for communication scenarios involving transmitter and receiver users from different linguistic backgrounds and facilitates the conversion of multimodal data from speech to text. The model structure of robust semantic communications for S2TT is shown in Fig. 1. From the figure, when the system input is clear speech, the deep semantic encoder compresses the speech sequence, $\boldsymbol{s}$, and extracts the deep semantic features, $\boldsymbol{f}$. Note that the deep semantic encoder directly learns the textual semantics in the target language. $\boldsymbol{f}$ is mapped to the symbols, $\boldsymbol{x}$, after passing through the NN-enabled channel encoder prior to transmission over the wireless channel. Then $\boldsymbol{x}$ can be expressed according to $\boldsymbol{s}$ as follows,
$\boldsymbol{x}={\mathfrak{T}}_{\mathrm{C}}({\mathfrak{T}}_{\mathrm{S}}(\boldsymbol{s}))\;\;\;\mathrm{w}.\mathrm{r}.\mathrm{t}.\;\;\;\boldsymbol{\alpha},$ (1)
where ${\mathfrak{T}}_{\mathrm{S}}(\cdot)$ and ${\mathfrak{T}}_{\mathrm{C}}(\cdot)$ are the deep semantic encoder and the channel encoder, respectively. $\boldsymbol{\alpha}$ is all trainable NN parameters of ${\mathfrak{T}}_{\mathrm{S}}(\cdot)$ and ${\mathfrak{T}}_{\mathrm{C}}(\cdot)$.
The encoded symbols, $\boldsymbol{x}$, are affected by the channel fading and channel noise after passing through the wireless channel layer, and the received symbols, $\boldsymbol{y}$, can be denoted as
@@
At the receiver, the symbols, $\boldsymbol{y}$, are recovered to the deep semantic features, $\widehat{\boldsymbol{f}}$, by the channel decoder. Moreover, the target text, $\widehat{\widetilde{\boldsymbol{t}}}$, is obtained by feeding $\widehat{\boldsymbol{f}}$ into the deep semantic decoder, which can be modelled as
$\widehat{\widetilde{\boldsymbol{t}}}=\mathfrak{T}_{\mathrm{S}}^{-1}(\mathfrak{T}_{\mathrm{C}}^{-1}\left(\boldsymbol{y}\right))\;\;\;\;\mathrm{w}.\mathrm{r}.\mathrm{t}.\;\;\;\boldsymbol{\beta},$ (3)
where $\mathfrak{T}_{\mathrm{C}}^{-1}(\cdot)$ and $\mathfrak{T}_{\mathrm{S}}^{-1}(\cdot)$ denotes the channel decoder and the deep semantic decoder, respectively. $\boldsymbol{\beta}$ represents all trainable NN parameters of $\mathfrak{T}_{\mathrm{C}}^{-1}(\cdot)$ and $\mathfrak{T}_{\mathrm{S}}^{-1}(\cdot)$.
@@ II-B Corrupted Speech Input
In practical communication systems, the integrity of input speech is highly susceptible to perturbations induced by the surrounding environments or the unstable network connection, thereby resulting in the potential degradation of original speech. In this work, our endeavours towards semantic communications for S2TT extend to the corrupted speech input, wherein some semantic information within the speech is inaccessible due to the introduction of semantic impairments. As shown in Fig. 1, the corrupted speech, $\overline{\boldsymbol{s}}$, contains limited semantic information. Particularly, the deep semantic compensator tasks $\overline{\boldsymbol{s}}$ as input and generates the compensated deep semantic features, $\widetilde{\boldsymbol{f}}$, which predicts the lost information and extracts textual semantic features in the target language simultaneously, written as
$\widetilde{\boldsymbol{f}}={\mathfrak{T}}_{\mathrm{SC}}(\overline{\boldsymbol{s}})\;\;\;\mathrm{w}.\mathrm{r}.\mathrm{t}.\;\;\;\boldsymbol{\gamma},$ (4)
where ${\mathfrak{T}}_{\mathrm{SC}}(\cdot)$ indicates the deep semantic compensator and $\boldsymbol{\gamma}$ is the corresponding trainable NN parameters.
Furthermore, a semantic impairment probe, $\boldsymbol{c}$, containing an index vector with position information of corrupted semantic features, is attained and transmitted to the receiver over a reliable channel. The motivation behind the semantic impairment probe is to strengthen the semantic fidelity of the recovered deep semantic features by further reducing the semantic ambiguity caused by corrupted semantic features. It is noteworthy that the channel encoder, channel decoder, and deep semantic decoder trained in the context of clear speech input are shared in communication scenarios with corrupted speech input to generate the target text.
@@ III Robust Semantic Communications for Speech Transmission
To address the preceding challenges, we adopt a two-stage training scheme. Particularly, a semantic transmission paradigm for S2TT based on clear speech input, named DeepSC-S2T, is first proposed. Then, a dual-compensator mechanism is carried out to enhance robust semantic communications, named Ross-S2T, which utilizes a GAN-enabled deep semantic compensator and a semantic impairment probe-aided compensator to acquire as accurate deep semantic features as possible at the receiver.
@@ III-A DeepSC-S2T
The proposed Ross-S2T is shown in Fig 2. From the figure, at the first training stage, the convolutional neural network (CNN) module condenses the clear speech and the transformer module further extracts the features, $\boldsymbol{F}$, before passing through the dense layer-enabled channel encoder to attain symbols, $\boldsymbol{X}$. The dense layer constructs the channel decoder to process the receiver symbols, $\boldsymbol{Y}$, and the transformer-enabled deep semantic decoder is leveraged to produce multiple target text sequences, $\widehat{\widetilde{\boldsymbol{T}}}$. To boost the efficient semantic transmission for serving the S2TT task, the label-smoothing regularization-aided cross-entropy (LSR-CE) is adopted as the E2E loss function to train the DeepSC-S2T, which is expressed as
@@
where $\kappa\in\left[0,1\right]$ is a hyperparameter that signifies the confidence level associated with the predicted tokens in $\widehat{\widetilde{\boldsymbol{T}}}$ matching the true tokens in the accurate text sequence in the target language, $\widetilde{\boldsymbol{T}}$. The trainable parameters $\boldsymbol{\theta}=(\boldsymbol{\alpha},\;\boldsymbol{\beta})$. $\widetilde{L}$ represents the number of tokens in $\widetilde{\boldsymbol{T}}$. ${\mathcal{L}}_{\mathrm{CE}}$ is the CE loss, denoted as
${\mathcal{L}}_{\mathrm{CE}}(\widetilde{\boldsymbol{T}},\;\widehat{\widetilde{\boldsymbol{T}}};\;\boldsymbol{\theta})=-\sum_{\widetilde{l}=1}^{\widetilde{L}}p({\widetilde{t}}_{\widetilde{l}})\log p({\widehat{\widetilde{t}}}_{\widetilde{l}}),$ (6)
where $p({\widetilde{t}}_{\widetilde{l}})$ and $p({\widehat{\widetilde{t}}}_{\widetilde{l}})$ are the true and predicted probabilities of token ${\widetilde{t}}_{\widetilde{l}}$ and ${\widehat{\widetilde{t}}}_{\widetilde{l}}$, respectively. The token $w_{e}$ belongs to a vocabulary group containing $E$ tokens and $w_{e}\neq{\widetilde{t}}_{\widetilde{l}}$. Therefore, $f_{w}\left(w_{e}\right)$ describes the confidence level associated with ${\widehat{\widetilde{t}}}_{\widetilde{l}}$ matching $w_{e}$ instead of ${\widetilde{t}}_{\widetilde{l}}$, which can be written as
@@
Figure 2: The model structure of proposed Ross-S2T for robust semantic communications with clear and corrupted speech inputs.
The intuition behind loss ${\mathcal{L}}_{\mathrm{LSR-CE}}$ is to introduce a level of confusion in predicting the target text, which enables the training uncertainty but ultimately improves the prediction accuracy in the testing stage. Algorithm 1 illustrates the algorithm for training the DeepSC-S2T.

### v3

@@ Abstract
In this paper, we propose a robust semantic communication system for speech transmission, named Ross-S2T, by delivering the essential semantic information. Specifically, we consider the speech-to-text translation (S2TT) as the transmission goal. First, a new deep semantic encoder is developed to convert speech in the source language to textual features associated with the target language, facilitating the end-to-end semantic exchange to perform the S2TT task and reducing the transmission data without performance degradation. To mitigate semantic impairments inherent in the corrupted speech, a novel generative adversarial network (GAN)-enabled deep semantic compensator is established to estimate the lost semantic information within the speech and extract deep semantic features simultaneously, which enables robust semantic transmission for corrupted speech. Furthermore, a semantic probe-aided compensator is devised to enhance the semantic fidelity of recovered semantic features and improve the understandability of the target text. According to simulation results, the proposed Ross-S2T exhibits superior S2TT performance compared to conventional approaches and high robustness against semantic impairments.
@@ I Introduction
According to the three-level communication architecture proposed by Shannon and Weaver [3], semantic communications are the second level of communications that prioritize conveying the underlying meaning by representing the input message with minimal ambiguity through semantics, which overcomes the limitation of conventional communications to process data at the bit level and drives the evolution of intelligent communications. However, there is no consensus on the definition of semantics, hindering the representation of semantic information with a rigorous formula. Thanks to the thriving of artificial intelligence (AI) in diverse areas, deep learning (DL)-enabled semantic communication paradigm breaks the bottleneck of a mathematical theory to quantify the semantics and has shown its great potential to learn semantic information through sophisticated neural networks (NNs). DL-enabled semantic communications have experienced unprecedented developments due to the ubiquity of intelligent mobile devices and the booming demand for semantic-driven data transmission.
@@
Particularly, the pioneering work on DL-enabled semantic communications, named DeepSC [4], has been proposed to recover the accurate text. A variant of DeepSC, named R-DeepSC [5], has been devised to eliminate semantic noise and facilitate robust text transmission. In [6], Weng et al. introduced a semantic communication system for speech transmission, named DeepSC-S, to extract and transmit global semantic features. Inspired by DeepSC-S, a deep speech semantic transmission scheme has been developed in [7] by adopting a flexible rate-distortion trade-off to achieve end-to-end (E2E) optimization. Huang et al. [8] considered a semantic communication system for image transmission. Jiang et al. [9] presented a video semantic transmission paradigm.
@@
Furthermore, Xie et al. [10] established a task-oriented semantic communication system for machine translation and visual question-answering tasks by fusing textual and visual semantic features. Weng et al. [11] designed a speech semantic transmission scheme, named DeepSC-ST, to perform speech recognition and synthesis tasks, and further explored the speech-to-text translation (S2TT) and speech-to-speech translation (S2ST) tasks in [12] by incorporating a machine translation module into DeepSC-ST. In [13], Xu et al. proposed reinforcement learning-enabled semantic communications for scene classification in unmanned aerial systems. Zhang et al. investigated a semantic communication system for extended reality (XR) tasks by transmitting highly compressed semantic information to reduce network traffic. In [14], a unified multimodal multi-task semantic communication architecture, named U-DeepSC, has been developed by sharing trainable parameters amongst various tasks to reduce the redundancy of semantic features and accelerate the inference process.
@@
In this paper, a robust semantic communication system for speech transmission, named Ross-S2T, is proposed. We argue that existing works on task-oriented semantic communications for speech only extract textual semantic features constrained to the source language, i.e., shallow semantic features, encouraging the exploration of deep semantic features spanning various languages. Moreover, we investigate the intractable semantic impairments. In this context, the S2TT task is considered in semantic transmission scenarios with corrupted speech input. The contributions of this paper are summarized as follows:
  • A semantic communication system for S2TT, named DeepSC-S2T, is developed to obtain the deep semantic features associated with the target language by leveraging a semantic extractor and a novel semantic converter.
@@
  • According to our comprehensive literature review, speech semantic impairments are not investigated in semantic communications. We propose a generative adversarial network (GAN)-enabled deep semantic compensator to estimate the damaged information in the corrupted speech and generate deep semantic features simultaneously.
@@ II System Model
This section introduces robust semantic communication systems for speech transmission and considers S2TT the transmission goal. The system aims to address two primary challenges. The first challenge is to deliver E2E semantic exchange and achieve efficient transmission from speech in the source language to text in the target language. The second is to devise a semantic impairment suppression mechanism to reduce semantic impairments within the corrupted speech. To this end, the novel deep semantic coding mechanism is established to facilitate speech transmission for S2TT, and the deep semantic compensator is first developed to compensate for the sophisticated semantic impairments.
@@ II-A Clear Speech Input
The designed system is tailored for communication scenarios involving transmitter and receiver users from different linguistic backgrounds and facilitates the conversion of multimodal data from speech to text. The framework of robust semantic communications for S2TT is shown in Fig. 1. From the figure, when the system input is clear speech, the deep semantic encoder compresses the speech sequence, $\boldsymbol{s}$, and produces the deep semantic features, $\boldsymbol{f}$. Note that the existing work on semantic communications for S2TT [12] only extracts the shallow semantic features related to the source language. It also generates intermediate source text at the receiver before producing the final target text, which hinders E2E training and requires additional computational resources for the machine translation module. To resolve this issue, we devise a novel semantic converter to transform shallow semantic features into deep semantic features, as shown in Fig. 2. From the figure, $\boldsymbol{f}$ can be expressed according to $\boldsymbol{s}$ as follows,
$\boldsymbol{f}={\mathfrak{T}}_{\mathrm{SC}}({\mathfrak{T}}_{\mathrm{SE}}(\boldsymbol{s}))\;\;\;\mathrm{w}.\mathrm{r}.\mathrm{t}.\;\;\;\boldsymbol{\alpha},$ (1)
where ${\mathfrak{T}}_{\mathrm{SE}}(\cdot)$ and ${\mathfrak{T}}_{\mathrm{SC}}(\cdot)$ are the semantic extractor and the semantic converter, respectively. ${\mathfrak{T}}_{\mathrm{DS}}=({\mathfrak{T}}_{\mathrm{SE}},\;{\mathfrak{T}}_{\mathrm{SC}})$. $\boldsymbol{\alpha}$ is all trainable NN parameters of ${\mathfrak{T}}_{\mathrm{DS}}$.
Figure 1: Model structure of robust semantic communications for speech-to-text translation.
Figure 2: The proposed deep semantic encoder.
The channel encoder maps $\boldsymbol{f}$ to the symbols, $\boldsymbol{x}$, before transmission over the wireless channel, denoted as,
$\boldsymbol{x}={\mathfrak{T}}_{\mathrm{C}}(\boldsymbol{f})\;\;\;\mathrm{w}.\mathrm{r}.\mathrm{t}.\;\;\;\boldsymbol{\beta},$ (2)
where ${\mathfrak{T}}_{\mathrm{C}}(\cdot)$ is the channel encoder and $\boldsymbol{\beta}$ is its NN parameters.
The encoded symbols, $\boldsymbol{x}$, are affected by the channel fading and channel noise after passing through the wireless channel layer, and the received symbols, $\boldsymbol{y}$, can be written as
@@
At the receiver, the target text, $\widehat{\widetilde{\boldsymbol{t}}}$, is obtained by feeding the recovered features, $\widehat{\boldsymbol{f}}$, into the deep semantic decoder, which can be modelled as
$\widehat{\widetilde{\boldsymbol{t}}}=\mathfrak{T}_{\mathrm{DS}}^{-1}(\mathfrak{T}_{\mathrm{C}}^{-1}\left(\boldsymbol{y}\right))\;\;\;\;\mathrm{w}.\mathrm{r}.\mathrm{t}.\;\;\;\boldsymbol{\rho},$ (4)
where $\mathfrak{T}_{\mathrm{C}}^{-1}(\cdot)$ and $\mathfrak{T}_{\mathrm{DS}}^{-1}(\cdot)$ denotes the channel decoder and the deep semantic decoder, respectively. $\boldsymbol{\rho}$ represents their trainable NN parameters.
@@ II-B Corrupted Speech Input
In practical communication systems, the integrity of input speech is highly susceptible to perturbations induced by surrounding environments or unstable network connections, resulting in the potential degradation of original speech. In this work, our endeavors towards semantic communications for S2TT extend to the corrupted speech input, wherein some semantic information within the speech is inaccessible due to the introduction of semantic impairments. Semantic impairments refer to external interference that damages the integrity of speech information. As shown in Fig. 1, the corrupted speech, $\overline{\boldsymbol{s}}$, contains limited semantic information. Particularly, the deep semantic compensator tasks $\overline{\boldsymbol{s}}$ as input and generates the compensated deep semantic features, $\widetilde{\boldsymbol{f}}$, which predicts the lost information and extracts textual semantic features in the target language simultaneously, written as
$\widetilde{\boldsymbol{f}}={\mathfrak{T}}_{\mathrm{DC}}(\overline{\boldsymbol{s}})\;\;\;\mathrm{w}.\mathrm{r}.\mathrm{t}.\;\;\;\boldsymbol{\delta},$ (5)
where ${\mathfrak{T}}_{\mathrm{DC}}(\cdot)$ indicates the deep semantic compensator and $\boldsymbol{\delta}$ is the corresponding trainable NN parameters.
Furthermore, a semantic impairment probe, $\boldsymbol{c}$, containing an index vector with position information of the corrupted semantic features, is attained and transmitted to the receiver over a reliable channel. The motivation behind the semantic impairment probe is to strengthen the semantic fidelity of the recovered deep semantic features by further reducing the semantic ambiguity caused by corrupted semantic features.
@@ III Robust Semantic Communications for Speech Transmission
To address the preceding challenges, we adopt a two-stage training scheme. Particularly, a semantic transmission paradigm for S2TT based on clear speech input, named DeepSC-S2T, is first proposed. Then, a dual-compensator mechanism is developed to enhance robust semantic communications, named Ross-S2T, which utilizes a GAN-enabled deep semantic compensator and a semantic impairment probe-aided compensator to acquire as accurate deep semantic features as possible at the receiver.
@@ III-A DeepSC-S2T
The proposed Ross-S2T is shown in Fig. 3. From the figure, at the first training stage, the convolutional neural network (CNN) module condenses the clear speech and the transformer module further extracts the features, $\boldsymbol{F}$, before passing through the dense layer-enabled channel encoder to attain symbols, $\boldsymbol{X}$. The dense layer constructs the channel decoder to process the receiver symbols, $\boldsymbol{Y}$, and the transformer-enabled deep semantic decoder is leveraged to produce multiple target text sequences, $\widehat{\widetilde{\boldsymbol{T}}}$. To boost the efficient semantic transmission for serving the S2TT task, the label-smoothing regularization-aided cross-entropy (LSR-CE) is adopted as the E2E loss function to train the DeepSC-S2T, which is expressed as
@@
where $\kappa\in\left[0,1\right]$ is a hyperparameter that signifies the confidence level associated with the predicted tokens in $\widehat{\widetilde{\boldsymbol{T}}}$ matching the true tokens in $\widetilde{\boldsymbol{T}}$. Note that $\widetilde{\boldsymbol{T}}$ is the accurate text sequence in the target language, $\widetilde{\boldsymbol{T}}$. The trainable parameters $\boldsymbol{\theta}=(\boldsymbol{\alpha},\;\boldsymbol{\beta},\;\boldsymbol{\rho})$. $\widetilde{L}$ represents the number of tokens in $\widetilde{\boldsymbol{T}}$. ${\mathcal{L}}_{\mathrm{CE}}$ is the CE loss. The token $w_{e}$ belongs to a vocabulary group containing $E$ tokens and $w_{e}\neq{\widetilde{t}}_{\widetilde{l}}$. Additionally, $f_{w}\left(w_{e}\right)$ describes the confidence level associated with ${\widehat{\widetilde{t}}}_{\widetilde{l}}$ matching $w_{e}$ instead of ${\widetilde{t}}_{\widetilde{l}}$, which can be written as
@@
The intuition behind loss ${\mathcal{L}}_{\mathrm{LSR-CE}}$ is to introduce a level of confusion in predicting the target text, which enables training uncertainty but ultimately improves prediction accuracy in the testing stage.
@@
Figure 3: Model structure of Ross-S2T for robust semantic communications with clear and corrupted speech inputs.